Find the number of atoms in 7.85 g of aluminum


1
Expert's answer
2021-05-31T02:03:31-0400

MM(Al) = 26.98 g/mol

n(Al) "= \frac{7.85}{26.98}= 0.29 \;mol"

A mole of Al contains 6.02×1023 atoms.

N(atoms)"= 0.29 \times 6.2 \times 10^{23}=1.8 \times 10^{23} \; atoms"